OpenStudy (anonymous):

In order to solve the equation -3x - 6 = 5, the first step is to:

OpenStudy (anonymous):

-3x - 6 = 5 transfer -6 to right side..

OpenStudy (anonymous):

makin co-efficient of x positive

OpenStudy (anonymous):

u can do this by transferring 5 to left and -3x to right

OpenStudy (anonymous):

\[-3x=5+6\]

OpenStudy (anonymous):

-3x - 6 = 5 -3x - 6 +6= 5+6 -3x=11 x= ?

OpenStudy (anonymous):

So I first start off by adding 6 to both sides?

OpenStudy (anonymous):

\[x=5+6/-3\]

OpenStudy (anonymous):